OpenStudy (nicolehagberg):

a model car has a scale of 1:25. the length of the model when built is 9.6 in, what is the length of the actual car

jimthompson5910 (jim_thompson5910):

`1:25` means you multiply the value 9.6 by 25 to get 9.6*25 = 240 so the actual length of the car is 240 inches convert this to feet: 240 inches * (1 ft/12 inches) = 240/12 = 20 feet the length of the car is 20 feet

OpenStudy (zzr0ck3r):

lol not even gonna try and make them work for it

OpenStudy (acespeedfighter):

███████████▓▓▓▓▓▓▓▓▒░░░░░▒▒░░░░░░░▓█████ ██████████▓▓▓▓▓▓▓▓▒░░░░░▒▒▒░░░░░░░░▓████ █████████▓▓▓▓▓▓▓▓▒░░░░░░▒▒▒░░░░░░░░░▓███ ████████▓▓▓▓▓▓▓▓▒░░░░░░░▒▒▒░░░░░░░░░░███ ███████▓▓▓▓▓▓▓▓▒░░▒▓░░░░░░░░░░░░░░░░░███ ██████▓▓▓▓▓▓▓▓▒░▓████░░░░░▒▓░░░░░░░░░███ █████▓▒▓▓▓▓▓▒░▒█████▓░░░░▓██▓░░░░░░░▒███ ████▓▒▓▒▒▒░░▒███████░░░░▒████░░░░░░░░███ ███▓▒▒▒░░▒▓████████▒░░░░▓████▒░░░░░░▒███ ██▓▒▒░░▒██████████▓░░░░░▓█████░░░░░░░███ ██▓▒░░███████████▓░░░░░░▒█████▓░░░░░░███ ██▓▒░▒██████████▓▒▒▒░░░░░██████▒░░░░░▓██ ██▓▒░░▒███████▓▒▒▒▒▒░░░░░▓██████▓░░░░▒██ ███▒░░░░▒▒▒▒▒▒▒▒▒▒▒▒░░░░░░███████▓░░░▓██ ███▓░░░░░▒▒▒▓▓▒▒▒▒░░░░░░░░░██████▓░░░███ ████▓▒▒▒▒▓▓▓▓▓▓▒▒▓██▒░░░░░░░▓███▓░░░░███ ██████████▓▓▓▓▒▒█████▓░░░░░░░░░░░░░░████ █████████▓▓▓▓▒▒░▓█▓▓██░░░░░░░░░░░░░█████ ███████▓▓▓▓▓▒▒▒░░░░░░▒░░░░░░░░░░░░██████ ██████▓▓▓▓▓▓▒▒░░░░░░░░░░░░░░░░▒▓████████ ██████▓▓▓▓▓▒▒▒░░░░░░░░░░░░░░░▓██████████ ██████▓▓▓▓▒▒██████▒░░░░░░░░░▓███████████ ██████▓▓▓▒▒█████████▒░░░░░░▓████████████ ██████▓▓▒▒███████████░░░░░▒█████████████ ██████▓▓░████████████░░░░▒██████████████ ██████▓░░████████████░░░░███████████████ ██████▓░▓███████████▒░░░████████████████ ██████▓░███████████▓░░░█████████████████ ██████▓░███████████░░░██████████████████ ██████▓▒██████████░░░███████████████████ ██████▒▒█████████▒░▓████████████████████ ██████░▒████████▓░██████████████████████ ██████░▓████████░███████████████████████ ██████░████████░▒███████████████████████ █████▓░███████▒░████████████████████████ █████▒░███████░▓████████████████████████ █████░▒██████░░█████████████████████████ █████░▒█████▓░██████████████████████████ █████░▓█████░▒██████████████████████████ █████░▓████▒░███████████████████████████ █████░▓███▓░▓███████████████████████████ ██████░▓▓▒░▓████████████████████████████ ███████▒░▒██████████████████████████████ ████████████████████████████████████████ ████████████████████████████████████████